OpenStudy (diyadiya):

Roots of the quadratic equation \( (a+b-2c)x^2-(2a-b+c)x + (a-2b+c) =0 \) are :

OpenStudy (anonymous):

what have you tried?

OpenStudy (diyadiya):

i tried to to -b/a and c/a :S

OpenStudy (diyadiya):

so what am i supposed to do?

OpenStudy (anonymous):

You have a+b and ab can't you find a and b ?

OpenStudy (diyadiya):

how?

OpenStudy (anonymous):

Find a-b and then eliminate.

OpenStudy (diyadiya):

where did a-b come from?

OpenStudy (anonymous):

It's a tedious process, I got a fairly critical form of r\mots: \[ \frac{1}{2} \left(\frac{(2 a-b+c)^2}{(a+b-2 c)^2}\pm \sqrt{\frac{9 b^2-22 b c+c (8 a+9 c)}{(a+b-2 c)^2}}\right) \]

OpenStudy (diyadiya):

Oh ok thanks !

OpenStudy (diyadiya):

Actually the answer was none of the options